OpenStudy (anonymous):

If 4 is a root of x^2+kx-48=0, find the other root and the value of k.

OpenStudy (cwrw238):

(x - 4)(x + 12) = x^2 + 8x - 48 so compare this to x^2 + kx - 48 = 0 to find k

OpenStudy (cwrw238):

understand?

OpenStudy (anonymous):

But where did you find the values x+12 and k?

OpenStudy (cwrw238):

4 is a given root so one factor of the trinomial x^2 + kx - 48 must be (x - 4) and in order to get the -48 the other factor must be (x + 12)

OpenStudy (cwrw238):

expanding (x-4)(x+12) gives x^2 + 12x - 4x - 48 = x^2 + 8x - 48

OpenStudy (cwrw238):

now if we compare this with the original equation: x^2 + 8x - 48 = 0 x^2 + kx - 48 = 0 comparing 8x and kx: k must be 8

OpenStudy (cwrw238):

and other root is -12

OpenStudy (cwrw238):

plz ask if you need more explanation

OpenStudy (anonymous):

Oh. Is there another method?

OpenStudy (cwrw238):

there could be but the above method is ok - i was really asking if you understood it

OpenStudy (anonymous):

I did. But it was more speculation than mathematical method.

OpenStudy (cwrw238):

not speculation at all

OpenStudy (anonymous):

To factor, the factors of the constant term have to add to the linear term. Given that x=4 is a root, (x-4) has to be a factor. Therefore, in order to get the -48 constant term, (x+12) needs to be the other factor. This will give k = 8 when you evaluate the product.